OpenStudy (anonymous):

Add 3x+9/2x+6 + 8x+12/x^2+6x+9 ??

OpenStudy (anonymous):

https://banff.owschools.com/media/g_alg02_2013/11/t12.gif For clarification

OpenStudy (solomonzelman):

I would expect that what you attached is the actual problem, even without attaching it :) factor out the denominators, to help to determine the common denominator.

OpenStudy (anonymous):

 it needs to be in this form

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{3x+9}{2x+6}+\frac{8x+12}{x^2+6x+9} }\) \(\LARGE\color{blue}{ \frac{3(x+3)}{2(x+3)}+\frac{4(2x+3)}{(x+3)(x+3)} }\) \(\LARGE\color{blue}{ \frac{3}{2}+\frac{4(2x+3)}{(x+3)(x+3)} }\) \(\LARGE\color{blue}{ \frac{3\color{red}{(x+3)(x+3)}}{2\color{red}{(x+3)(x+3)}}+\frac{4(2x+3)\color{red}{\times 2} }{(x+3)(x+3)\times 2} }\) \(\LARGE\color{blue}{ \frac{3(x+3)(x+3)}{2(x+3)(x+3)}+\frac{8(2x+3) }{2(x+3)(x+3)} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{3(x^2+6x+9)}{2(x+3)(x+3)}+\frac{8(2x+3) }{2(x+3)(x+3)} }\) \(\LARGE\color{blue}{ \frac{3x^2+18x+27}{2(x+3)(x+3)}+\frac{16x+24 }{2(x+3)(x+3)} }\) \(\LARGE\color{blue}{ \frac{3x^2+18x+27+16x+24}{2(x+3)(x+3)} }\) \(\LARGE\color{blue}{ \frac{3x^2+34x+51}{2(x+3)(x+3)} }\)